a piece of wire 10 m long is to be cut into two pieces one of them is 2/3 as long as the other, how should the wire be cut ?
Answer by checkley77(12844)

X+2X/3=10
(3X+2X)/3=10
5X=3810
5X=30
X=30/5
X=6 M. IS THE LONGER PIECE.
10-6=4 M. IS THE SHORTER PIECE.
PROOF:
6+2*6/3=10
6+12/2=10
6+4=10
10=10